Operations Report – January 29th

Crew 220 (MAU) Operations Report 29JAN2020

SOL 10

Name of person filing report: Lee Roberts, Crew Engineer

Non-nominal systems: N/A

Generator: Operating nominally.

Hours run: Not run overnight

From what time last night: N/A

To what time this morning: N/A

List of any additional daytime hours when the generator was run: 2 hours

Solar- SOC% (Before the generator is run at night): N/A

Diesel Reading – N/A

Station Propane Reading- 100% + Backup (1.8 cylinders)

Ethanol Free Gasoline- N/A

Water (loft tank) (gal) – N/A

Water Meter (units)- N/A

Water (static tank) (gal) – N/A

Static to loft Pump used – N/A

Water in GreenHab (gal): N/A

Water in Science Dome (gal): N/A

Toilet Tank Emptied – no

Deimos rover used: N/A


Beginning Charge:

End Charge:

Currently Charging:

Sojourner rover used: Assigned to Director


Beginning Charge:

End Charge:

Currently Charging:

Spirit Rover used – Used for exploration EVA

Starting Hours: 128.3

Beginning Charge: 100%

Ending Hours: 128.6

Ending Charge: 63%

Opportunity Rover used: Still in workshop


Beginning Charge:

Ending Charge:

Currently Charging: Yes

Curiosity Rover used: Used for exploration EVA

Starting Hours: 134.3

Beginning Charge: 100%

Ending Hours: 134.6

Ending Charge: 62%

Notes on Rovers: Opportunity and Deimos off-site for maintenance.

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): No

Reason for use:

Oil Added? No

# hours the ATV’s were used today:

Notes on ATVs:

Habcar used and why, where? Used to town to get water.

CrewCar used and why, where? Not used.

General Notes and Comments:

Summary of the internet: N/A

Summary of Suit and Radios: Nominal

Summary of Hab: The crew spent several hours this morning cleaning the Station and trying to mitigate mud intrusion on the floors. Tarps were placed on the floor in two of the rooms to reinforce the floor surfaces.

During external inspection of the Station it was found that several supporting poles for the Station walls were dislodged during the recent windstorm, compromising the structural integrity. These poles were reaffixed, and the walls appear to be structurally sound.

In addition to replenishing the propane supply, we also installed a charger to maintain our backup battery at a steady charge, and we are instituting a power management plan to reduce propane consumption on the generator.

Summary of Science Dome operations: Temperatures are hovering around 5°C during the day.

Summary of RAM operations: Nominal

Summary of any observatory issues: Nominal

Summary of health and safety issues: Nominal

Questions, concerns, and requests to Mission Support: None

Copyright © The Mars Society. All rights reserved. | Main Site